OpenStudy (anonymous):

alfie and bon travelled at rates of 50kph and 60kph respectively.Alfie traveled 2 hrs longer than bon and went 50 km farther. what is the distance traveled by alfie?

OpenStudy (anonymous):

if alphie traveled 2 hrs at 50 kph then he traveled 100 kilometers after bon stopped but was only 50 ahead. that means at the time bon stopped, bon was 50 ahead of alfie. Since alfie would lose 60-50 =10 kil per hour they traveled 5 hours before bon stopped which means alfie traveled 5 + 2 = 7 hours. 7 hours at 50 kph = 350 kilometers traveled by alfie and 60*5 = 300 killometers traveled by bon. Double checking we see a diference of 2 hours and 50 kil so we are correct. the answer is that alfie traveled 350 killometers.